# Incompressible Limit of Compressible Ideal MHD Flows inside a Perfectly Conducting Wall

### v4

@@ 3.3.1 The $\mathcal{T}^{\alpha}$-differentiated equations
$\varepsilon^{8}\mathcal{T}^{\alpha}\text{ with }\langle\alpha\rangle=8\text{ and }\alpha_{0}=0,~{}~{}\varepsilon^{8}\partial_{t}^{8},~{}~{}\varepsilon^{k-1}\partial_{t}^{k}~{}(4\leq k\leq 7).$
@@ 3.3.2 Tangential energy estimate with full spatial derivatives
$\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}(B\cdot\nabla B)\cdot(\varepsilon^{8}\overline{\partial}{}^{8}u)\,\mathrm{d}x-\int_{\Omega}\varepsilon^{2k}\overline{\partial}{}^{8}(\nabla(p+|B|^{2}/2))\cdot(\varepsilon^{2k}\overline{\partial}{}^{4+k}u)\,\mathrm{d}x.$ (3.66)
@@
$-\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}p\cdot(\varepsilon^{8}\overline{\partial}{}^{8}(\varepsilon^{2}D_{t}p))\,\mathrm{d}x$
$=$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|\varepsilon^{8}\overline{\partial}{}^{8}(\varepsilon p)\right|^{2}\,\mathrm{d}x\underbrace{-\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}p\cdot(\varepsilon^{10}[\overline{\partial}{}^{8},u\cdot\nabla]p)\,\mathrm{d}x}_{:=\mathcal{J}_{4}}.$ (3.70)
@@
The last term $\mathcal{J}_{6}$ is again directly controlled: $\mathcal{J}_{6}\leq P(E(t))$. The first term cancels with $\mathcal{K}_{2}$. For the second term, we need to further analyze the case when $j=3$. Thanks to the continuity equation, we have $-\partial_{3}u_{3}=\varepsilon^{2}D_{t}p+\overline{\nabla}\cdot\bar{u}$ and thus
$-\sum_{i,j=1}^{3}\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}B_{i}\cdot(\varepsilon^{8}\overline{\partial}{}B_{i}\overline{\partial}{}^{7}\partial_{3}u_{3})\,\mathrm{d}x=\sum_{i=1}^{3}\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}B_{i}\cdot\left(\varepsilon^{8}\overline{\partial}{}B_{i}\overline{\partial}{}^{7}(\varepsilon^{2}D_{t}p+\overline{\nabla}\cdot\bar{u})\right)\,\mathrm{d}x\lesssim P(E(t)).$ (3.74)
When $j=1,2$, the second term is again directly controlled by $P(E(t))$ because all the derivatives are tangential.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|\varepsilon^{8}\overline{\partial}{}^{8}(\varepsilon p,u,B)\right\|_{0}^{2}\leq P(E(t)).$ (3.77)
@@ 3.3.3 Tangential energy estimate with full time derivatives
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\sum_{k=4}^{7}\left\|\varepsilon^{k-1}\partial_{t}^{k}(u,B,\varepsilon p)\right\|_{0}^{2}+\left\|\varepsilon^{8}\partial_{t}^{8}(u,B,\varepsilon p)\right\|_{0}^{2}\right)\leq P(E(t)).$ (3.82)
@@
$-\int_{\Omega}\varepsilon^{8}\partial_{t}^{8}p\left(\varepsilon^{10}[\partial_{t}^{8},u\cdot\nabla]p\right)\,\mathrm{d}x=-\int_{\Omega}\varepsilon^{9}\partial_{t}^{8}p\left(\varepsilon^{9}[\partial_{t}^{8},u\cdot\nabla]p\right)\,\mathrm{d}x$ (3.83)
$=$ $-\int_{\Omega}\varepsilon^{9}\partial_{t}^{8}p\left(\varepsilon^{9}(\partial_{t}^{8}u_{j}\partial_{j}p+8\partial_{t}u_{j}\partial_{j}\partial_{t}^{8}p)\right)\,\mathrm{d}x$
$-\int_{\Omega}(\varepsilon^{9}\partial_{t}^{8}p)\left(\sum_{k=2}^{7}\binom{8}{k}(\varepsilon^{k-1}\partial_{t}^{k}u_{j})(\varepsilon^{10-k}u_{j}\partial_{j}\partial_{t}^{8-k}p)\right)\,\mathrm{d}x\lesssim E(t).$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|\varepsilon^{8}\partial_{t}^{8}(u,B,\varepsilon p)\right\|_{0}^{2}\leq P(E(t)).$ (3.84)
@@
$-\int_{\Omega}\varepsilon^{6}\partial_{t}^{7}p\left(\varepsilon^{8}[\partial_{t}^{7},u\cdot\nabla]p\right)\,\mathrm{d}x=-\int_{\Omega}\varepsilon^{7}\partial_{t}^{7}p\left(\varepsilon^{7}[\partial_{t}^{7},u\cdot\nabla]p\right)\,\mathrm{d}x$ (3.85)
$=$ $-\int_{\Omega}\varepsilon^{7}\partial_{t}^{7}p\left(\varepsilon^{7}(\partial_{t}^{7}u_{j}\partial_{j}p+7\partial_{t}u_{j}\partial_{j}\partial_{t}^{6}p)\right)\,\mathrm{d}x$
$-\int_{\Omega}(\varepsilon^{7}\partial_{t}^{7}p)\left(\sum_{k=2}^{6}\binom{7}{k}(\varepsilon^{k-1}\partial_{t}^{k}u_{j})(\varepsilon^{8-k}u_{j}\partial_{j}\partial_{t}^{8-k}p)\right)\,\mathrm{d}x\lesssim E(t).$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|\varepsilon^{6}\partial_{t}^{7}(\varepsilon p,u,B)\right\|_{0}^{2}\leq P(E(t)).$ (3.86)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\sum_{k=4}^{7}\left\|\varepsilon^{k-1}\partial_{t}^{k}(u,B,\varepsilon p)\right\|_{0}^{2}\leq P(E(t)).$ (3.87)
@@ 3.3.4 Space-time mixed tangential derivatives
$\frac{\mathrm{d}}{\mathrm{d}t}\bigg{(}$ $\sum_{\langle\alpha\rangle=8,\alpha_{0}<8}\left\|\varepsilon^{8}\mathcal{T}^{\alpha}(u,B,p)\right\|_{0}^{2}+\left\|\varepsilon^{8}\partial_{t}^{8}(u,B,\varepsilon p)\right\|_{0}^{2}$ (3.88)
$+\sum_{l=0}^{3}\sum_{\langle\alpha\rangle=2l,\alpha_{0}<2l}\left\|\varepsilon^{3+l}\mathcal{T}^{\alpha}\partial_{t}^{4-l}(u,B,p)\right\|_{0}^{2}+\sum_{l=0}^{3}\left\|\varepsilon^{3+l}\partial_{t}^{4+l}(u,B,\varepsilon p)\right\|_{0}^{2}\bigg{)}\leq P(E(t)).$
@@ 3.4 Uniform estimates
Since $D_{t}S=0$, we can easily prove the estimates for $S$ by directly commuting $D_{t}$ with *$\partial_{*}^{\alpha}$ and corresponding weights of Mach number. The proof does not involve any boundary term because we do not integrate by parts, so we omit the details.
@@ 3.5 The case of 2D MHD
$\|X\|_{s}^{2}\lesssim\|X\|_{0}^{2}+\|\nabla\cdot X\|_{s-1}^{2}+\|\nabla^{\perp}\cdot X\|_{s-1}^{2}+|\overline{\partial}{}_{1}X\cdot N|^{2}_{s-\frac{3}{2}},$ (3.92)
@@
$\mathcal{K}^{\prime\prime}:=\int_{\Omega}\partial^{3}(\nabla^{\perp}\cdot B)~{}\partial^{3}\nabla^{\perp}\cdot(\varepsilon^{2}BD_{t}p)\,\mathrm{d}x,$ (3.94)
@@
$\varepsilon^{2}(B\cdot\nabla^{\perp})\partial^{3}D_{t}p=\varepsilon^{2}(-B_{1}\partial^{3}\partial_{2}D_{t}p+B_{2}\partial^{3}\partial_{1}D_{t}p)$
@@
$\varepsilon^{2}(-B_{1}\partial^{3}\partial_{2}D_{t}p+B_{2}\partial^{3}\partial_{1}D_{t}p)$
$=$ $~{}\varepsilon^{2}\left(-B_{1}\partial^{3}D_{t}(\partial_{2}p)+B_{2}\partial^{3}D_{t}(\partial_{1}p)\right)-\varepsilon^{2}\left(B_{1}\partial^{3}(\partial_{2}u_{j}\partial_{j}p)-B_{2}\partial^{3}(\partial_{1}u_{j}\partial_{j}p)\right),$
@@
$\varepsilon^{2}(-B_{1}\partial^{3}\partial_{2}D_{t}p+B_{2}\partial^{3}\partial_{1}D_{t}p)$
@@
$\mathcal{K}^{\prime\prime}\stackrel{{\scriptstyle L}}{{=}}$ $\int_{\Omega}\partial^{3}(\nabla^{\perp}\cdot B)\varepsilon^{2}\rho(B_{1}\partial^{3}D_{t}^{2}u_{2}-B_{2}\partial^{3}D_{t}^{2}u_{1})\,\mathrm{d}x-\int_{\Omega}\partial^{3}(\nabla^{\perp}\cdot B)~{}\varepsilon^{2}|B|^{2}\partial^{3}D_{t}(\nabla^{\perp}\cdot B)\,\mathrm{d}x$ (3.95)
=$\stackrel{{\scriptstyle L}}{{=}}$ $\int_{\Omega}\partial^{3}(\nabla^{\perp}\cdot B)\varepsilon^{2}\rho(B_{1}\partial^{3}D_{t}^{2}u_{2}-B_{2}\partial^{3}D_{t}^{2}u_{1})\,\mathrm{d}x$
$-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|\varepsilon|B|\partial^{3}(\nabla^{\perp}\cdot B)\right|^{2}\,\mathrm{d}x+\frac{1}{2}\int_{\Omega}(\nabla\cdot u)\left(\left|\varepsilon|B|\partial^{3}(\nabla^{\perp}\cdot B)\right|^{2}\right)\,\mathrm{d}x,$
where the last term on the right side can be directly controlled, and the first term requires the control of $\|\varepsilon^{2}D_{t}^{2}u\|_{3}$. The conclusion of Lemma 3.2 is then modified to be
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho\left|\partial^{3}(\nabla^{\perp}\cdot u)\right|^{2}+(1+\varepsilon^{2}|B|^{2})\left|\partial^{3}(\nabla^{\perp}\cdot B)\right|^{2}\,\mathrm{d}x\leq P(E_{4}(t))+E_{5}(t).$ (3.96)

### v6

@@ 3.3.1 The $\mathcal{T}^{\alpha}$-differentiated equations
$\varepsilon^{8}\mathcal{T}^{\alpha}\text{ with }\langle\alpha\rangle=8\text{ and }\alpha_{0}=0,~{}~{}\varepsilon^{8}\partial_{t}^{8},~{}~{}\varepsilon^{k-1}\partial_{t}^{k}~{}(4\leq k\leq 7)$
@@ 3.3.2 Tangential energy estimate with full spatial derivatives
$\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}(B\cdot\nabla B)\cdot(\varepsilon^{8}\overline{\partial}{}^{8}u)\,\mathrm{d}x-\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}(\nabla(p+|B|^{2}/2))\cdot(\varepsilon^{8}\overline{\partial}{}^{4}u)\,\mathrm{d}x.$ (3.67)
@@
$-\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}p\cdot(\varepsilon^{8}\overline{\partial}{}^{8}({\mathcal{F}_{p}}D_{t}p))\,\mathrm{d}x$
=${\stackrel{{\scriptstyle L}}{{=}}}$ $-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|\varepsilon^{8}{\mathcal{F}_{p}^{\frac{1}{2}}}\overline{\partial}{}^{8}p\right|^{2}\,\mathrm{d}x\underbrace{-\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}p\cdot(\varepsilon^{8}{\mathcal{F}_{p}}[\overline{\partial}{}^{8},u\cdot\nabla]p)\,\mathrm{d}x}_{:=\mathcal{J}_{4}}.$ (3.71)
@@
The last term $\mathcal{J}_{6}$ is again directly controlled: $\mathcal{J}_{6}\leq P(E(t))$. The first term cancels with $\mathcal{K}_{2}$. For the second term, we need to further analyze the case when $j=3$. Thanks to the continuity equation, we have $-\partial_{3}u_{3}={\mathcal{F}_{p}}D_{t}p+\overline{\nabla}\cdot\bar{u}$ and thus
$-\sum_{i,j=1}^{3}\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}B_{i}\cdot(\varepsilon^{8}\overline{\partial}{}B_{i}\overline{\partial}{}^{7}\partial_{3}u_{3})\,\mathrm{d}x=\sum_{i=1}^{3}\int_{\Omega}\varepsilon^{8}\overline{\partial}{}^{8}B_{i}\cdot\left(\varepsilon^{8}\overline{\partial}{}B_{i}\overline{\partial}{}^{7}({\mathcal{F}_{p}}D_{t}p+\overline{\nabla}\cdot\bar{u})\right)\,\mathrm{d}x\lesssim P(E(t)).$ (3.75)
When $j=1,2$, the second term is again directly controlled by $P(E(t))$ because all derivatives are tangential.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|\varepsilon^{8}\overline{\partial}{}^{8}({\mathcal{F}_{p}^{\frac{1}{2}}}p,u,B)\right\|_{0}^{2}\leq P(E(t)).$ (3.78)
@@ 3.3.3 Tangential energy estimate with full time derivatives
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\sum_{k=4}^{7}\left\|\varepsilon^{k-1}\partial_{t}^{k}(u,B,{\mathcal{F}_{p}^{\frac{1}{2}}}p)\right\|_{0}^{2}+\left\|\varepsilon^{8}\partial_{t}^{8}(u,B,\varepsilon p)\right\|_{0}^{2}\right)\leq P(E(t)).$ (3.83)
@@
$-\int_{\Omega}\varepsilon^{8}\partial_{t}^{8}p\left(\varepsilon^{8}{\mathcal{F}_{p}}[\partial_{t}^{8},u\cdot\nabla]p\right)\,\mathrm{d}x=-\int_{\Omega}\varepsilon^{8}{\mathcal{F}_{p}^{\frac{1}{2}}}\partial_{t}^{8}p\left(\varepsilon^{8}{\mathcal{F}_{p}^{\frac{1}{2}}}[\partial_{t}^{8},u\cdot\nabla]p\right)\,\mathrm{d}x$ (3.84)
$=$ $-\int_{\Omega}\varepsilon^{8}{\mathcal{F}_{p}^{\frac{1}{2}}}\partial_{t}^{8}p\left(\varepsilon^{8}{\mathcal{F}_{p}^{\frac{1}{2}}}(\partial_{t}^{8}u_{j}\partial_{j}p+8\partial_{t}u_{j}\partial_{j}\partial_{t}^{8}p)\right)\,\mathrm{d}x$
$-\int_{\Omega}(\varepsilon^{8}{\mathcal{F}_{p}^{\frac{1}{2}}}\partial_{t}^{8}p)\left(\sum_{k=2}^{7}\binom{8}{k}(\varepsilon^{k-1}\partial_{t}^{k}u_{j})(\varepsilon^{9-k}{\mathcal{F}_{p}^{\frac{1}{2}}}u_{j}\partial_{j}\partial_{t}^{8-k}p)\right)\,\mathrm{d}x\lesssim E(t).$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|\varepsilon^{8}\partial_{t}^{8}(u,B,{\mathcal{F}_{p}^{\frac{1}{2}}}p)\right\|_{0}^{2}\leq P(E(t)).$ (3.85)
@@
$-\int_{\Omega}\varepsilon^{6}\partial_{t}^{7}p\left(\varepsilon^{6}{\mathcal{F}_{p}}[\partial_{t}^{7},u\cdot\nabla]p\right)\,\mathrm{d}x=-\int_{\Omega}\varepsilon^{6}{\mathcal{F}_{p}^{\frac{1}{2}}}\partial_{t}^{6}{\mathcal{F}_{p}^{\frac{1}{2}}}p\left(\varepsilon^{7}[\partial_{t}^{7},u\cdot\nabla]p\right)\,\mathrm{d}x$ (3.86)
$=$ $-\int_{\Omega}\varepsilon^{6}{\mathcal{F}_{p}^{\frac{1}{2}}}\partial_{t}^{7}p\left(\varepsilon^{6}{\mathcal{F}_{p}^{\frac{1}{2}}}(\partial_{t}^{7}u_{j}\partial_{j}p+7\partial_{t}u_{j}\partial_{j}\partial_{t}^{6}p)\right)\,\mathrm{d}x$
$-\int_{\Omega}(\varepsilon^{6}{\mathcal{F}_{p}^{\frac{1}{2}}}\partial_{t}^{7}p)\left(\sum_{k=2}^{6}\binom{7}{k}(\varepsilon^{k-1}\partial_{t}^{k}u_{j})(\varepsilon^{7-k}{\mathcal{F}_{p}^{\frac{1}{2}}}u_{j}\partial_{j}\partial_{t}^{8-k}p)\right)\,\mathrm{d}x\lesssim E(t).$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|\varepsilon^{6}\partial_{t}^{7}(u,B,{\mathcal{F}_{p}^{\frac{1}{2}}}p)\right\|_{0}^{2}\leq P(E(t)).$ (3.87)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\sum_{k=4}^{7}\left\|\varepsilon^{k-1}\partial_{t}^{k}(u,B,{\mathcal{F}_{p}^{\frac{1}{2}}}p)\right\|_{0}^{2}\leq P(E(t)).$ (3.88)
@@ 3.3.4 Space-time mixed tangential derivatives
$\frac{\mathrm{d}}{\mathrm{d}t}\bigg{(}$ $\sum_{\langle\alpha\rangle=8,\alpha_{0}<8}\left\|\varepsilon^{8}\mathcal{T}^{\alpha}(u,B,p)\right\|_{0}^{2}+\left\|\varepsilon^{8}\partial_{t}^{8}(u,B,{\mathcal{F}_{p}^{\frac{1}{2}}}p)\right\|_{0}^{2}$ (3.89)
$+\sum_{l=0}^{3}\sum_{\langle\alpha\rangle=2l,\alpha_{0}<2l}\left\|\varepsilon^{3+l}\mathcal{T}^{\alpha}\partial_{t}^{4-l}(u,B,p)\right\|_{0}^{2}+\sum_{l=0}^{3}\left\|\varepsilon^{3+l}\partial_{t}^{4+l}(u,B,{\mathcal{F}_{p}^{\frac{1}{2}}}p)\right\|_{0}^{2}\bigg{)}\leq P(E(t)).$
@@ 3.4 Uniform estimates
Since $D_{t}S=0$, we can easily prove the estimates for $S$ by directly commuting $D_{t}$ with $\partial_{*}^{\alpha}$ and corresponding weights of Mach number. The proof does not involve any boundary term because we do not integrate by parts, so we omit the details.
@@ 3.5 The case of 2D MHD
$\|X\|_{s}^{2}\lesssim\|X\|_{0}^{2}+\|\nabla\cdot X\|_{s-1}^{2}+\|\nabla^{\perp}\cdot X\|_{s-1}^{2}+|X\cdot N|^{2}_{s-\frac{1}{2}},$ (3.93)
@@
$\mathcal{K}^{\prime\prime}:=\int_{\Omega}\partial^{3}(\nabla^{\perp}\cdot B)~{}\partial^{3}\nabla^{\perp}\cdot({\mathcal{F}_{p}}BD_{t}p)\,\mathrm{d}x,$ (3.95)
@@
${\mathcal{F}_{p}}(B\cdot\nabla^{\perp})\partial^{3}D_{t}p={\mathcal{F}_{p}}(-B_{1}\partial^{3}\partial_{2}D_{t}p+B_{2}\partial^{3}\partial_{1}D_{t}p)$
@@
${\mathcal{F}_{p}}(-B_{1}\partial^{3}\partial_{2}D_{t}p+B_{2}\partial^{3}\partial_{1}D_{t}p)$
$=$ $~{}{\mathcal{F}_{p}}\left(-B_{1}\partial^{3}D_{t}(\partial_{2}p)+B_{2}\partial^{3}D_{t}(\partial_{1}p)\right)-{\mathcal{F}_{p}}\left(B_{1}\partial^{3}(\partial_{2}u_{j}\partial_{j}p)-B_{2}\partial^{3}(\partial_{1}u_{j}\partial_{j}p)\right),$
@@
${\mathcal{F}_{p}}(-B_{1}\partial^{3}\partial_{2}D_{t}p+B_{2}\partial^{3}\partial_{1}D_{t}p)$
@@
$\mathcal{K}^{\prime\prime}\stackrel{{\scriptstyle L}}{{=}}$ $\int_{\Omega}\partial^{3}(\nabla^{\perp}\cdot B){\mathcal{F}_{p}}\rho(B_{1}\partial^{3}D_{t}^{2}u_{2}-B_{2}\partial^{3}D_{t}^{2}u_{1})\,\mathrm{d}x-\int_{\Omega}\partial^{3}(\nabla^{\perp}\cdot B)~{}{\mathcal{F}_{p}}|B|^{2}\partial^{3}D_{t}(\nabla^{\perp}\cdot B)\,\mathrm{d}x$ (3.96)
=$\stackrel{{\scriptstyle L}}{{=}}$ $\int_{\Omega}\partial^{3}(\nabla^{\perp}\cdot B){\mathcal{F}_{p}}\rho(B_{1}\partial^{3}D_{t}^{2}u_{2}-B_{2}\partial^{3}D_{t}^{2}u_{1})\,\mathrm{d}x$
$-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\left|{\mathcal{F}_{p}^{\frac{1}{2}}}|B|\partial^{3}(\nabla^{\perp}\cdot B)\right|^{2}\,\mathrm{d}x+\frac{1}{2}\int_{\Omega}(\nabla\cdot u)\left(\left|{\mathcal{F}_{p}^{\frac{1}{2}}}|B|\partial^{3}(\nabla^{\perp}\cdot B)\right|^{2}\right)\,\mathrm{d}x,$
where the last term on the right side can be directly controlled, and the first term requires the control of $\|{\mathcal{F}_{p}}D_{t}^{2}u\|_{3}$. The conclusion of Lemma 3.2 is then modified to be
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\Omega}\rho\left|\partial^{3}(\nabla^{\perp}\cdot u)\right|^{2}+(1+{\mathcal{F}_{p}}|B|^{2})\left|\partial^{3}(\nabla^{\perp}\cdot B)\right|^{2}\,\mathrm{d}x\leq P(E_{4}(t))+E_{5}(t).$ (3.97)